OpenStudy (anonymous):

If y varies inversely as x, and y = 5 as x = -2, find y for the x-value of 3.

OpenStudy (shinalcantara):

If y varies inversely as x, then you'll have it as \[y = \frac{1}{x}k\] where k is the constant of proportionality ----------------------------------- Given that y=5 when x=-2, substitute to the above function to find 'k' \[y = \frac{1}{x}k\] \[5 = \frac{ 1 }{ -2 } k \] Cross multiply -2 \[5(-2) = k\] \[k = -10\] ----------------------------------- Using the same relation but when x=3, y=? \[y = \frac{1}{x}k\] Substitute x=3 and k=-10 \[y = \frac{1}{x}k\] \[y = \frac{1}{3}(-10)\] \[y=?\]